OpenStudy (anonymous):

If A = (7 – 11 + 8)^131 and B = (–7 + 11 – 8)^131 , then what is the value of (7 – 13)^(A+B) ? Please explain.

OpenStudy (anonymous):

\[A =4^{131}\] \[B =-4^{131}\] \[A +B =4^{131}-4^{131}=0\] \[\left( 7-13 \right)^{A +B}=\left( -6 \right)^{0}=1\]

OpenStudy (anonymous):

Thank you so much!